OpenStudy (anonymous):

solve by elmination method 8x -9y = 24.5 7y -2x = 8.5

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

(8x-9y)+4(7y-2x)=24.5+4*8.5 19y=58.5 y=3.079 and x=6.527

OpenStudy (anonymous):

say the answer should be an ordered pair?

OpenStudy (anonymous):

yes...(6.527,3.079)

OpenStudy (anonymous):

it shows answer when i submitted it it said it should have been (2.5, -0.5)

OpenStudy (anonymous):

this is wrong because (2.5,-.5) does not satisfy ur second equation...